OpenStudy (credmond):

Find the indicated power using De Moivre's Theorem. (Express your fully simplified answer in the form a + bi.) (3(sqrt)3 + 3i)^-5

OpenStudy (anonymous):

@credmond Do you know what De'Moivre's Theorem is and what it's used for?

OpenStudy (credmond):

mhm!

OpenStudy (credmond):

do you need it?

OpenStudy (dumbcow):

is there an "i" as part of the exponent?

OpenStudy (credmond):

oh shoot there shouldn't be. sorry that was a typo! just raised to the -5 power!

OpenStudy (dumbcow):

start by rewriting in polar form \[\rightarrow [6(\cos \frac{\pi}{6} +i \sin \frac{\pi}{6})]^{-5i}\] ... oh shoot i had it all worked out even with the "i" haha

OpenStudy (dumbcow):

so anyway ignore the "i" then applying the theorem \[\rightarrow 6^{-5} (\cos \frac{-5\pi}{6} +i \sin \frac{-5\pi}{6})\]

OpenStudy (dumbcow):

\[= -\frac{\sqrt{3}}{2}6^{-5} -\frac{1}{2}6^{-5} i\]

OpenStudy (credmond):

Thank you so much!! Sorry about the typo! haha @dumbcow

OpenStudy (dumbcow):

yw

OpenStudy (kenljw):

(3(sqrt)3 + 3i)^-5 3^-5 ((sqrt)3 + i)^-5 3^-5 4^-5 (cos(PI/6) + i sin(PI/6))^-5 (cos(-5 PI/6) + i sin(-5 PI/6)) 12^-5

OpenStudy (credmond):

thank you @KenLJW !